CH/Meet/93/06

Minutes

Charissa Group Meeting

Meeting held on: 17 November 1993

Meeting held at: Birmingham

Present:

BRF(Chair),WNC,NMC,MF,GJG,SJAH,RAH,JSL,ACM,VP, WDMR,GT,RPW,DLW plus P.H. Regan (for Items 1, 2 and 3.)

Apologies:  RAC,CDJ,JK
Distribution: Those above plus DDW, all by e-mail/file-server.

Minutes

1. Minutes of meeting of 9 September: accepted

2. Matters Arising/ Actions:

• Item 10: Beamline at ANU — BRF had approached H.Price and the NPC. Formal agreement had been obtained to use ex–NSF equipment and possibly a Birmingham pump. The installation would be studied by the support group, along with the planned beamline for Strasbourg. Contact with

ANU to be maintained.  JSL

• Agenda The present agenda was deliberately kept brief because the main items demanded considerable discussion. It was agreed that the omission of other regular items was regretted but essential.

3. De-briefing on ANU Running:

The inaugural running period in Australia had been successfully completed.

BRF extended thanks on behalf of everyone to the advance group (Mike Leddy, Garry Tungate and Steve Hall) and the “basic presence” group (Steve Hall, Roger Ward and Chris Jones) for their successful efforts. Special thanks went to Steve Hall for his long stint, and special commendation to Paddy Regan (ANU) for his valuable assistance.

BRF had asked in advance of the meeting for people to be ready to suggest items for attention before the next running period. DLW tabled a list of comments on these matters, prepared by WDMR and himself during the recent running.

It was agreed to work through the tabled list, and actions were assigned to 26 agreed items. A document detailing these actions should be prepared. WNC

read document ch/doc/93/14 to check actions on individuals Available on the file-server: /charissa/document/doc9314/ps

4. Discussion on Grant Submission:

BRF had circulated a draft copy of the Charissa document that would be submitted as an Annex to each group’s grant request.

The draft document was discussed in detail: report, future plans, resources requested.

DLW pointed out that grant requests should make clear that requests were costed according to academic years.

The costings of new developments and the breakdown of these between institutions were agreed. The manpower to be requested under each institution was also agreed.

It was agreed that a meeting to discuss radioactive beam initiatives was re-

quired.  BRF

It was agreed that each institution would attend to its own travel requests independently from the Charissa annex. The costing scheme employed for technical Eurogam travel was noted, viz. mid-week European fares, hotel room with breakfast, plus £25/day subsistence. An estimate of Charissa travel costs was derived as follows: 2 trips to the ANU at £2K/trip plus European travel and conferences, giving £6K/yr per person all in.

5. Any Other Business:

BRF reported that he had discussed the developments in data acquisition with VP, and that it was recommended that the readout into VME would be put in place for March 1994. This was agreed.

6. Date of Next Meeting:

Next Meeting: 11.30 a.m., Wednesday, 5 January 1993 in Birmingham.

Lunch arrangements to be made.  BRF

W.N. Catford, Secretary. November 30, 1993 (released December 1, 1993)